# Stochastically Structured Reservoir Computers for Financial and Economic System Identification

### v1

@@ 4 Algorithms
Compute economy-sized SVD $USV=A$;
Set $s=\min\{m,n\}$;
Set $r=\sum_{j=1}^{s}H_{\varepsilon}(s_{j}(A))$;
Set $U_{\delta}=\sum_{j=1}^{r}U\hat{e}_{j,s}\hat{e}_{j,s}^{\ast}$;
Set $T_{\delta}=\sum_{j=1}^{r}(\hat{e}_{j,s}^{\ast}S\hat{e}_{j,s})^{-1}\hat{e}_{j,s}\hat{e}_{j,s}^{\ast}$;
Set $V_{\delta}=\sum_{j=1}^{r}\hat{e}_{j,s}\hat{e}_{j,s}^{\ast}V$;
Set $\hat{A}=U_{\delta}^{\ast}A$;
Set $\hat{y}=U_{\delta}^{\ast}y$;
Set $x_{0}=V_{\delta}^{\ast}T_{\delta}\hat{y}$;
@@
       Set $x=\mathbf{0}_{n,1}$;
       Set $A_{0}=\sum_{j=1}^{N_{0}}\hat{A}\hat{e}_{\sigma(j),n}\hat{e}_{j,N_{0}}^{\ast}$;
       Solve $c=\arg\min_{\tilde{c}\in(\mathbb{R}^{+}_{0})^{N_{0}}}\|A_{0}\tilde{c}-\hat{Y}\hat{e}_{j,p}\|$;
@@
Algorithm 1 SLRSolver: Sparse linear least squares solver algorithm
Data: $\Sigma_{T}=\{x(t)\}_{t=1}^{T}\subset\mathbb{R}^{n},\mathbb{B}_{\mathcal{S}}\subset\mathbb{R}^{m\times n}$
Result: $\hat{W},\tilde{\eth}_{p,r}$
  • 1:
    Set a tensor order value $p$.
  • 2:
    Compute the reduced embedding ð~$\tilde{\eth}_{p,r}$ applying the map $R_{p}(n)$ determined by Lemma 2.
  • 3:
    Compute matrices:
    $\mathbf{X}_{0}(T)$ $:=R_{p}(n)\begin{bmatrix}|&&|\\ \tilde{\eth}_{p}\left(\mathbf{x}(0)\right)&\ldots&\tilde{\eth}_{p}\left(\mathbf{x}(T-1)\right)\\ |&&|\end{bmatrix}$
[…]
  • 4:
    Set:
    $X_{0}$ =[||$\ vec(S_{1}\mathbf{X}_{0}(T))&\cdots&vec(S_{q}\mathbf{X}_{0}(T))\\ |&&|\end{bmatrix}$
    $X_{1}$ $=vec(\mathbf{X}_{1}(T))$
    for $\mathbb{B}_{\mathcal{S}}:=\{S_{1},\ldots,S_{q}\}$
  • 5:
    Approximately solve:
    $\begin{bmatrix}X_{0}^{\top}X_{0}\\ C\end{bmatrix}\mathbf{a}=\begin{bmatrix}X_{0}^{\top}X_{1}\\ \mathbf{1}_{p}\end{bmatrix}$
    using Algorithm 2
  • 6:
    Set:
    $\hat{W}:=\sum_{j=1}^{q}\mathbf{a}[j]S_{j}$
  • 7:
    Set $\hat{W}:=\sum_{j=1}^{m}c_{j}\hat{X}_{j}$.
return $\hat{W},\tilde{\eth}_{p,r}$
Algorithm 2 SSRC Model: SSRC model identification
@@ 4.1 Interpretation
From a systems-theoretic perspective, the proposed algorithmic framework reveals a subtle yet powerful idea: change within a system does not necessarily require external shocks, but can instead emerge from the way the system internally represents and responds to its own evolving state. The embedding ð~$\tilde{\eth}_{p,r}(x)$, shaped by structure-preserving constraints and stochastic compression, serves not only as a computational mechanism but also as a formal reflection of the system’s internal logic.
In this sense, the dynamics modeled by equations such as (1) capture more than just numerical evolution, they encode the possibility that adaptation and feedback may arise from internal symmetries and symbolic representations. This offers a conceptual bridge between control, cognition, and structural self-awareness, and opens the door to new interpretations of autonomy and regulation in complex socio-economic systems.
@@ 5 Computational simulations
The dynamic models considered in this section are determined by particular representations of the switched control systems (1), (2).
5.1 Resource Competition Among Agents: A Dynamic and Behavioral Approach
@@
This simulation introduces a dynamic behavioral model of resource competition between multiple agents who interact strategically and adaptively to maximize their participation benefits. Rather than advancing normative claims regarding the desirability of maintaining a given level of participation, the analysis focuses on the evolving trajectories of resource allocation and the emergent patterns over time. The model considered in this study is given by
$\mathbf{p}(t+1)=W\,\tilde{\eth}_{2}(\mathbf{p}(t)),$ (5)
where ð~$\tilde{\eth}_{2}(\mathbf{p}(t))$ denotes a structured embedding of reduced second-degree monomials. The structural configuration of the maps under consideration allows us to interpret equation (5) as a nonlinear closed-loop control system of the form:
$\mathbf{p}(t+1)=\frac{1}{3}W_{1}\,\mathbf{p}(t)+\frac{1}{3}W_{2}\,u(\mathbf{p}(t))+\frac{1}{3}\,\mathbf{w}_{3},$ (6)
@@
where $u(\mathbf{p}(t))$ is a stochastic vector composed of reduced monomials of degree two, $\mathbf{w}_{3}$ is a constant stochastic vector, and $W_{1},W_{2}$ are stochastic matrices of appropriate dimensions that represent state transitions and feedback interactions, respectively.
The structured output coupling matrix obtained from the identification of a resource competition model of the form (5), using a set of synthetic signals provided in Vides (2023), is shown in Figure 1, alongside the corresponding relational graph.
@@
Figure 1: Structured coupling matrix (left) and associated relational graph (right) for the resource competition model.
@@
The results of the signal identification process based on models (5) and (6) are presented in Figure 2, showing both short-term and long-term resource distribution estimates.
@@
Figure 2: Estimated short-term resource distribution (left) and long-term distribution under closed-loop dynamics (right).
The model demonstrates that, under certain conditions, competitive dynamics can lead to concentration processes in which resources initially distributed among $n$ agents progressively accumulate around a smaller subset. Such outcomes may arise both from active accumulation strategies and from the adaptive withdrawal of agents who choose to exit the competitive environment in favor of alternative domains.
@@
Due to their flexible structure, models of the form (6) can be applied across a variety of domains, including social, economic, financial, ecological, and symbolic systems. Although it does not impose a predefined causal structure, the results raise theoretical and empirical questions about the conditions under which resource concentration may be systematically associated with broader dynamics such as hierarchical stratification, exclusionary processes, or power asymmetries.
@@ 5.2 Regional Inflation Network Dynamics
Several studies have highlighed the significant influence of U.S. monetary policy on the monetary frameworks and inflation rates of other countries (Tenkovskaya, 2023; Azad and Serletis, 2022; Carella et al., 2024). Furthermore, the existence of cross-country inflation transmission dynamics has also been documented (Budová et al., 2023; Iraheta Bonilla et al., 2008; Liu et al., 2015).
Following the COVID-19 recession, many countries, including the United States, adopted expansionary monetary policies to mitigate the economic downturn, with the federal funds rate being a key tool (Feldkircher et al., 2021). According to Swanson (2024), changes in the federal funds rate significantly impact production and prices, highlighting short-term interest rates as the most powerful tool central banks have to influence the economy.
Beginning in 2022, the Federal Reserve adjusted its strategy in response to external factors and rising inflation, leading to higher federal rates and tighter financial conditions (Alekseievska et al., 2024). Most of the CAPARD countries implemented similar adjustments in response to regional inflation. These measures were followed by a gradual decline in inflation in both the US and the region, as shown in Figure 5.
Motivated by this context, we examine the interaction of inflation signals among a group of interconnected economies, including large ones such as the United States and China, and smaller regional economies such as those in the Central American region, Panama, and the Dominican Republic (CAPARD), along with the U.S. federal funds rate, included to capture the potential influence of U.S. monetary policy on countries inflation dynamics, over the period January 2020 to November 2024. For this purpose, we will consider structured switched models of the form:
$\mathbf{x}(t+1)=A_{\tau(t)}\mathbf{x}(t)$
Here, the structure of $A_{\tau(t)}$ is determined by some suitable economic interrelation constraints between the countries under consideration.
Our analysis yielded notable findings. First, we identified a network of interconnections between the inflationary dynamics of the countries in the Northern Triangle of Central America, inflation, and the U.S. federal funds rate, exhibiting varying lags in comparison to other countries in the region. The contributions to the inflationary states of these countries are depicted in Figure 4, through the structural identification of the matrix and its relational behavior, as shown in the empirical relational graph presented in Figure 3.
@@
$x(t+1)=W_{int}y(t)+W_{lat}z(t)$
Here $x(t+1)$ is explained by the dynamics of the real signals in $y(t)$. The model balances interpretability and predictive capability, offering accurate predictions while preserving context.
@@ 6 Conclusion and future work
This work presented a novel methodology for identifying and simulating financial and economic systems using stochastically structured reservoir computers (SSRCs). By integrating structure-preserving embeddings with coupling matrices constrained by relational graphs, the proposed approach enables interpretable and context-aware representations of dynamic systems. The identification process is grounded in a constrained optimization framework that ensures compliance with both stochastic and structural properties.
Through two empirical case studies, a dynamic model of resource competition among agents and a regional inflation network model, we demonstrated the capacity of SSRCs to capture complex behavioral patterns, simulate nonlinear feedback mechanisms, and uncover meaningful interdependencies in real-world economic signals. These results highlight the potential of structured reservoir computing not only as a modeling tool, but also as a lens to understand systems in which change emerges from the way agents interact with and represent their own environment.
Future research will focus on expanding the proposed methodology in three directions. First, by refining the latent signal modeling component to further enhance stability and contextual preservation in economic predictions. Second, by extending the structured embeddings to incorporate exogenous shocks and long-memory effects. Third, by developing adaptive learning schemes that allow online updating of coupling matrices as new data becomes available. These directions aim to consolidate SSRCs as a versatile and theoretically grounded tool for studying complex adaptive systems across domains.
@@
The programs and data that support the findings of this study will be openly available in the DyNet-CNBS repository, reference number Vides (2023), in due time.

### v3

@@ 4 Algorithms
Data: $\Sigma_{T}(x)=\{x(t)\}_{t=1}^{T}\subset\mathbb{R}^{n}$, $\Sigma_{T}(y)=\{y(t)\}_{t=1}^{T}\subset\mathbb{R}^{m},\mathbb{B}_{\mathcal{S}}(m,n)\subset\mathbb{R}^{m\times q}$
Result: $\hat{W},\eth_{p}$
  • 1:
    Choose or identify $s\in\{0,1\}^{p+1}$.
  • 2:
    Compute $\eth_{p}$ as either the reduced embedding $\tilde{\eth}_{s,p,r}$ or a function of it, using the map $R_{s,p}(n)$ from Lemma 2.
  • 3:
    Compute matrices:
    $X_{0}(T)$ $:=R_{p}(n)\begin{bmatrix}|&&|\\ \eth_{p}\left(x(0)\right)&\ldots&\eth_{p}\left(x(T-1)\right)\\ |&&|\end{bmatrix}$
    $X_{1}(T)$ $:=\begin{bmatrix}|&&|\\ y(0)&\ldots&y(T-1)\\ |&&|\end{bmatrix}$
  • 4:
    Set:
    $X_{0}$ $=\begin{bmatrix}|&&|\\ vec(S_{1}X_{0}(T))&\cdots&vec(S_{q}X_{0}(T))\\ |&&|\end{bmatrix}$
    $X_{1}$ $=vec(X_{1}(T))$
    for $\mathbb{B}_{\mathcal{S}}(m,n):=\{S_{1},\ldots,S_{q}\}$
  • 5:
    Approximately solve the structured nonnegative least squares problem:
    $\mathbf{a}:=\arg\min_{\hat{\mathbf{a}}\in(\mathbb{R}_{0}^{+})^{q}}\left\|\begin{bmatrix}X_{0}^{\top}X_{0}\\ C\end{bmatrix}\hat{\mathbf{a}}-\begin{bmatrix}X_{0}^{\top}X_{1}\\ \mathbf{1}_{p}\end{bmatrix}\right\|_{2}$
    using Algorithm 2
  • 6:
    Set:
    $\hat{W}:=\sum_{j=1}^{q}\mathbf{a}[j]S_{j}$
return $\hat{W},\eth_{p}$
Algorithm 1 SSRC Model: SSRC model identification
@@
1exSet $x_{0}=\arg\min_{\tilde{c}\in(\mathbb{R}^{+}_{0})^{n}}\|A\tilde{c}-y\|_{2}$;
@@
   Set $x=\mathbf{0}_{n}$;
   Set $A_{0}=\sum_{j=1}^{N_{0}}A\hat{e}_{\sigma(j),n}\hat{e}_{j,N_{0}}^{\ast}$;
   Solve $c=\arg\min_{\tilde{c}\in(\mathbb{R}^{+}_{0})^{N_{0}}}\|A_{0}\tilde{c}-\hat{Y}\hat{e}_{j,p}\|_{2}$;
@@
Algorithm 2 SSSLSSolver: Sparse stochastically structured linear least squares solver algorithm
@@ 4.1 Interpretation
From a control and optimization standpoint, the proposed framework formalizes systemic adaptation through structure-constrained embeddings and graph-restricted coupling matrices. The stochastic embedding $\tilde{\eth}_{p,r}(x)$ acts as a dimensional and structural regularizer, ensuring that the reconstructed representation remains consistent with the probabilistic and relational constraints of the system.
Accordingly, the dynamics in (1) perform more than state prediction: the embedding encodes feasible transformations within the stochastic manifold, while the optimization stage identifies the coupling matrix $W$ inside the admissible structural subspace $(\mathrm{span}\,\mathbb{B}_{\mathcal{S}}(m,n))\cap\mathbb{S}_{m,n}(\mathbb{R})$. This joint mechanism promotes interpretability and stability in the identified model.
Thus, the SSRC algorithm operates as a structure-preserving solver that unifies identification, constraint enforcement, and adaptive regularization, providing a transparent link between algorithmic operations and system-level behavior.
Remark. Models of the form (1) also allow the identification of autoregressive systems with control inputs. This is achieved by replacing $y(t)$ with $x(t+1)$ and redefining the embedding input as
$x(t)\;\mapsto\;\begin{bmatrix}x(t)\\[2.0pt] u(t)\end{bmatrix}$,
where $u(t)$ denotes an exogenous or control signal.
@@ 5 Computational simulations
The dynamic models considered in this section are determined by particular representations of the dynamic models (1).
5.1 Modeling Preference Evolution in Investment Coalitions with SSRCs
@@
This simulation introduces a stochastic dynamic model of resource allocation among agents who interact strategically and adaptively to maximize their participation benefits in the selection of an investment account within a given portfolio. Rather than advancing normative claims about the desirability of specific participation levels, the analysis focuses on the evolving trajectories of allocation and the emergent patterns over time. The model considered is given by
$p(t+1)=W\,\eth_{2}(p(t)),$ (3)
where $\eth_{2}(p(t))$ denotes a structured embedding of reduced second-degree monomials extracted from $\tilde{\eth}_{(1,1,0),2}(p(t))$, defined as
$\eth_{2}(p(t))=\begin{bmatrix}0.9\,p(t)\\ 0.1\,p(t)[1]\,p(t)[2]\\ 0.1\,p(t)[3]\,p(t)[5]\\ 0.1\,(1-p(t)[1]\,p(t)[2]-p(t)[3]\,p(t)[5])\end{bmatrix}$,
with $p(t)[j]$ denoting the probability of choosing account $j$ in the portfolio at time $t$.
The structural configuration of the maps in (3) allows us to interpret the system as a nonlinear dynamic model of the form:
$p(t+1)=0.9\,W_{1}\,p(t)+0.1\,W_{2}\,u(p(t)),$ (4)
@@
where $u(p(t))$ is a stochastic vector of reduced second-degree monomials, and $W_{1},W_{2}$ are stochastic matrices that represent state transitions and feedback interactions, respectively:
$W_{1}$ $:=\begin{bmatrix}1&0.45&0&0&0\\ 0&0&0.45&0&0\\ 0&0.55&0&0.45&0\\ 0&0&0.55&0&0\\ 0&0&0&0.55&1\end{bmatrix},\quad W_{2}:=\begin{bmatrix}0.8&0.1&0\\ 0&0&0\\ 0&0&0\\ 0&0&0\\ 0.2&0.9&1\end{bmatrix}$.
The structured output coupling matrix $\hat{W}$ identified from the model (3), using synthetic signals generated by iterating on it, is given by
$\hat{W}:=\begin{bmatrix}1&0.45&0&0&0&0.8&0.1&0\\ 0&0&0.45&0&0&0&0&0\\ 0&0.55&0&0.45&0&0&0&0\\ 0&0&0.55&0&0&0&0&0\\ 0&0&0&0.55&1&0.2&0.9&1\end{bmatrix}$,
with its sparsity pattern shown in Figure 1, together with the associated relational graph.
@@
Figure 1: Structured coupling matrix (left) and relational graph (right) of the investment preference evolution model.
@@
The results of the identification process based on models (3) and (4) are shown in Figure 2, illustrating the short- and long-term evolution of investment preference allocations.
@@
Figure 2: Estimated short-term evolution (left) and long-term evolution (right) of investment preference allocations.
This example shows that SSRCs can model stochastically embedded substochastic systems. By Lemma 1, any sub-vector extracted from tensor product blocks under a stochastic embedding defines a substochastic vector. This property will also be used in the next example to identify relational graphs among economic variables.
@@
Thanks to their flexible structure, models of the form (4) apply across social, economic, financial, ecological, and symbolic domains. Although they do not impose a predefined causal structure, the results raise questions about the conditions under which resource concentration may be linked to broader dynamics such as stratification, exclusion, or power asymmetries.
@@ 5.2 Regional Inflation Network Dynamics
Several studies have highlighted the significant influence of U.S. monetary policy on the monetary frameworks and inflation rates of other countries (Tenkovskaya2023; Azad2022; Carella2024). Furthermore, the existence of cross-country inflation transmission dynamics has also been documented (Budova2023; Iraheta2008; Liu2015).
Following the COVID-19 recession, many countries, including the United States, adopted expansionary monetary policies to mitigate the economic downturn, with the federal funds rate being a key tool (Feldkircher2021). According to Swanson2024, changes in the federal funds rate significantly impact production and prices, highlighting short-term interest rates as the most powerful tool central banks have to influence the economy.
Beginning in 2022, the Federal Reserve adjusted its strategy in response to external factors and rising inflation, leading to higher federal rates and tighter financial conditions (Alekseievska2024). Most of the CAPARD countries implemented similar adjustments in response to regional inflation. These measures were followed by a gradual decline in inflation in both the US and the region, as shown in Figure 5.
Motivated by this context, we examine the interaction of inflation signals among a group of interconnected economies, including large ones such as the United States and China, and smaller regional economies such as those in the Central American region, Panama, and the Dominican Republic (CAPARD), along with the U.S. federal funds rate, included to capture the potential influence of U.S. monetary policy on countries inflation dynamics, over the period January 2020 to November 2024. For this purpose, we will consider structured models of the form:
x(t+1)=Ax(t)x(t+1)=A~x(t)
Here, the structure of $A$ is determined by some suitable economic interrelation constraints between the countries under consideration.
Our analysis yielded notable findings. First, we identified a network of interconnections between the inflationary dynamics of the countries in the Northern Triangle of Central America inflation, and the U.S. federal funds rate, exhibiting varying lags in comparison to other countries in the region. The contributions to the inflationary states of these countries are depicted in Figure 4, through the structural identification of the matrix and its relational behavior, as shown in the empirical relational graph presented in Figure 3.
@@
$x(t+1)=W_{int}x(t)+W_{lat}z(t)$
Here $x(t+1)$ is explained by the dynamics of the real signals in $x(t)$. The model balances interpretability and predictive capability, offering accurate predictions while preserving context.
@@ 6 Conclusion and Future Work
This paper presented a methodology for identifying and simulating financial and economic systems through stochastically structured reservoir computers (SSRCs). By integrating structure-preserving embeddings with graph-constrained coupling matrices, the framework yields interpretable and structurally consistent representations of dynamic processes.
Applications to a nonlinear stochastic model and a regional inflation network demonstrated the capability of SSRCs to capture nonlinear behaviors, reveal interdependencies, and reproduce emergent patterns in economic data. These results confirm the efficiency of SSRCs as modeling tools and their potential for explaining systemic change driven by agent interactions.
Future work will focus on refining latent signal modeling for stability, extending embeddings to capture shocks and long-memory effects, and developing adaptive schemes for online learning. These advances aim to consolidate SSRCs as a robust framework for analyzing complex adaptive systems.
@@
The programs and data that support the findings of this study are openly available in the DyNet-CNBS repository, reference number FVides_DyNet.